Questions 3. Exchange Rate Effects on Investing. Explain how the appreciation of the Australian dollar against the U.S. dollar would affect the return to a U.S. firm that invested in an Australian money market security 4. Exchange Rate Effects on Borrowing. Explain how the appreciation of the Japanese yen against the U.S. dollar would affect the return to a U.S. firm that borrowed Japanese yen and used the proceeds for a U.S. project. 6. Bid/ask Spread. Utah Banks s bid price for Canadian dollars is S.7938 and its ask price is $.81. What is the bid/ask percentage spread? 9. Euro. Explain the foreign exchange situation for countries that use the euro when they engage in international trade among themselves. 10. Indirect Exchange Rate. If the direct exchange rate of the euro is $1.25, what is the euros indirect exchange rate? That is, what is the value of a dollar in euros? ● 36 Questions 11. Cross Exchange Rate. Assume Polands currency (the zloty) is worth S.17 and the Japanese yen is worth $.008. What is the cross rate of the zloty with respect to yen? That is, how many yen equal a zloty? 19, Foreign Stock Markets. Explain why firms may issue stock in foreign markets. Why might U.S. firms issue more stock in Europe since the conversion to the euro in 1999? 24. Interpreting Exchange Rate Quotations. Today you notice the following exchange rate quotations: $1 is equal to 3.00 Argentine pesos * 1 Argentine peso = 0.50 Canadian dollars You need to purchase 100,000 Canadian dollars with U.S. dollars. How many U.S. dollars will you need for your purchase?
Answer #1

Solution

Q3. To understand let's assume

  • 1USD = 1.5 AUD at the time of investment.
  • You invested 1USD = 1.5AUD in money market
  • Return on this investment is 0.2 AUD so Value of investment = 1.5+0.2 = 1.7AUD
  • Australian currency appreciated and new exchange rate is 1USD = 1AUD
  • If u convert this to USD then you will get 1.7 USD
Before appreciation After Appreciation
Exchange Rate : 1USD = 1.5AUD Exchange Rate : 1USD = 1AUD
Amount Invested: 1.5AUD Amount Invested: 1.5AUD
Return : 0.2 AUD Return : 0.2 AUD
Total: 1.7 AUD Total: 1.7 AUD
Convert in USD = 1.7/1.5 = 1.133 USD Convert in USD = 1.7/1 = 1.7 USD

So return will get increased due to appreciation of Australian currency

Question 4

Let's assume a scenario to understand this. The US firm has borrowed 100 yen and has 0% interest rate

Before appreciation After Appreciation
Exchange Rate : 1USD = 100 yen Exchange Rate : 1USD = 50 yen
Amount loaned: 100 yen Amount loaned: 100 yen
USD needed = 100 /100 = 1USD USD needed = 100 /50 =2 USD

The US firm has now have to pay more dollars if yen appreciates

In general, if you have taken a foreign currency loan and if foreign currency appreciates then you have to pay more

Question 6)

Bid = $0.7938 Ask = $0.81

Bid -ask spread = (Ask - Bid ) / Ask *100 =  (0.81 - 0.7938 ) / 0.81 *100 = 2%

Question 9)

Contries that uses Euro for trade can have have liability to pay in Euro.

Importers that imports from European Union or have liability in Euro

  • So if Euro appreciates against domestic currency then importer has to pay more and if Euro depreciates then they have better situation and have to pay less in domestic currency

Question 10)

1 Euro = $1.25 Direct quotation

1$ = 1/1.25 Euro = 0.8 Euro : Indirect quotation

Question 11)

1 Zloty = $.17

1 Yen = $0.008

1 $ = 1 / .17 Zloty = 5.882353 zloty

1$ = 1/0.008 Yen = 125 Yen

5.882353 zloty = 125 Yen .

1 zloty = 125 / 5.882353 yen = 21.25 Yen

Question 19)

  • Firm issues stocks in foreign market so that investors can invest in their home country and they will not get the exposure of exchange rate fluctuations as investment and return will be in domestic currency

Question 24)

1$ = 3 Argentine peso

1 Argentine peso = 0.5 Canadian dollar

3 Argentine peso = 1.5 Canadian dollar

1$ = 1.5 Canadian dollar

1 Canadian dollar need 1 /1.5 USD

100,000 canadian dollar will need 100,000 * 1/ 1.5 = 66,666.67 USD